The proportion of students that have taken the census is π = .7, what is the probability that a sample of size 50 yields a sample proportion of .75
Answer #1

Sol:

p=population proportion=0.7

p^=sample proportion=x/n=0.75

n=50

P(p^=0.75)

but z=p^-p/sqrt(p*(1-p)/n

z=0.75-0.7/sqrt((0.7)*(1-0.7)/50)

P(Z=0.7715)

from normal distribbution table

=0.7798

ANSWER:

0.7798